OpenStudy (anonymous):

What is the integration of dy/dt = (y/8) (6-y)

OpenStudy (anonymous):

first separate the variables dy/dt = (y/8)(6-y) dy = (y/8)(6-y) dt dy/(y/8)(6-y) = dt int( dy/(6y/8 - y^2/8) = int(dt) can you solve this?

OpenStudy (anonymous):

yeah I got up to that but I don't know how to solve after would that be ln of that value on the left side ?

OpenStudy (anonymous):

For this one I think you have to integrate by parts like so: \[\int\limits_{}^{}1/(y/8)(6-y)dy = \int\limits_{}^{}(A/(y/8) + B/(6-y))dy\] Have you learned this technique of integration yet?

OpenStudy (anonymous):

Oh I was doing it like this \[\ln \left| 6y-y^2 \right| = e^{x/8+C}\] And that's where I'm having trouble going on to the next step.

OpenStudy (anonymous):

e^ ln on the left sorry. And then I have 6y-y^2 = e^(x/8+C)

OpenStudy (anonymous):

I do not think that you integrated the left-hand term correctly. Try integrating by parts on the y term

OpenStudy (anonymous):

How do I integrate (6y-y^2)^-1 dy

OpenStudy (anonymous):

Here's how you integrate by parts: 1/(y/8)(6-y) = A/(y/8) + B/(6-y), where you solve for A and B First, find a common denomenator 1/(y/8)(6-y) = A(6-y)/(y/8)(6-y) + B(y/8)/(6-y)(y/8) Then get rid of the denominator by multiplying both sides by the denominator 1 = A(6-y) + B(y/8) Then set values for y so that you can solve for A and B individually let y = 0 1 = A(6-0) + B(0) A = 1/6 let y = 6 1 = A(0) + B(6/8) B = 8/6 or 4/3 plug in A and B int( (1/6)/(y/8) + (4/3)/(6-y))dy) this should be an integral you can solve. If you follow a textbook, they probably go into detail on how to solve these problems

OpenStudy (anonymous):

\[ \text{we have:}\quad \int{8\over y(6-y)}dy=\int dt\\ {\rm let}\quad {A\over y}+{B\over 6-y}={8\over y(6-y)}\\ \implies A(6-y)+By=8\\ {\rm let}\quad y=0\implies 6A=8\implies \boxed{A={4\over3}}\\ {\rm let}\quad y=6\implies 6B=8\implies \boxed{B={4\over3}}\\ \text{the integrand becomes:}\\ \int \left[{4/3\over y}+{4/3\over 6-y}\right]dy+C=t\\ \]this is now readily solvable

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

this is called integration by parts

OpenStudy (anonymous):

Uh I think this way is called something else. Integration by parts would use the formula u*v-integral v du.